# STRIVE: Structured Reasoning for Self-Improvement in Claim Verification

### v1

@@ 3.1 Task Formulation
Formally, given a claim $c$ and an evidence set =$\mathcal{E}=\{e_{1},e_{2},\dots,e_{n}\}$, where each $e_{i}$ is a descriptive piece of evidence (such as a sentence or paragraph), the goal is to obtain a model that can generate a reasoning chain $r$ that forms the intermediate reasoning steps. The final prediction $\hat{p}$ is made based on the reasoning chain, where $\hat{p}\in\{\textit{Supported},\textit{Refuted}\}$.
@@ 3.3.2 Reasoning Chain Generation and Selection
1:  $M^{*}\leftarrow\text{train}(M,\mathcal{D}_{h})$ // Structured Warm-up
2:  $\hat{r}_{i}\leftarrow M^{*}(T(c_{i},\mathcal{E}_{i})),\ \hat{p}_{i}\leftarrow J(\hat{r}_{i})\ \forall i\in[1,N]$ //Get reasoning chains and predicted labels
3:  $\mathcal{D}_{1}\leftarrow\{(c_{i},\mathcal{E}_{i},\hat{r}_{i})\mid\hat{p}_{i}=p_{i}\}$ //Select correct reasoning chains based on ground truth label
4:  $\hat{r}_{i}^{hint}\leftarrow M^{*}(T^{hint}(c_{i},\mathcal{E}_{i})),\ \hat{p}_{i}^{hint}\leftarrow J(\hat{r}_{i}^{hint}),$$\forall i\in[1,N],\text{ where }\hat{p}_{i}\neq p_{i}$ //Regenerate reasoning chains for incorrect predictions
5:  $\mathcal{D}_{2}\leftarrow\{(c_{i},\mathcal{E}_{i},\hat{r}_{i}^{hint})\mid\hat{p}_{i}\neq p_{i},\hat{p}_{i}^{hint}=p_{i}\}$//Select corrected reasoning chains
6:  $\mathcal{D}_{st}\leftarrow\{(c_{i},\mathcal{E}_{i},r_{i})\mid r_{i}\in(\mathcal{D}_{1}\cup\mathcal{D}_{2}),f(r_{i})=\text{True}\}$ //Select structurally valid chains
7:  $M_{st}\leftarrow\text{train}(M,\mathcal{D}_{st}\cup\mathcal{D}_{h})$ //Final Self-improvement training
@@
Format Checking: At this stage, both $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ contain reasoning chains that yield correct final predictions. To further ensure the quality of these reasoning chains, we apply a rule-based structural verification function $f$, where $f(\hat{r}_{i})=\textit{True}\text{ or }\textit{False}$, indicating whether $\hat{r}_{i}$ follow our predefined structure. Specifically, we enforce three criteria:
@@ A.4 STRIVE-Generated Reasoning Chains
Claim: The star of Forever Strong played a character that is based on the Hanna-Barbera show “Birdman and the Galaxy Trio”. Label: Supported Model Output: C1: The star of Forever Strong played a character that is based on the Hanna-Barbara show “Birdman and the Galaxy Trio”. Entity Resolution: Star of Forever Strong -> Gary Cole (from E1) Resolution Verification: E1 confirms Gary Cole was a star of Forever Strong. Entity Resolution: Character played by Gary Cole -> ? (from E2) Resolution Verification: E2 confirms Gary Cole voiced Harvey Birdman, but it doesn’t mention him playing a character in Forever Strong. Verification: E1 shows Gary Cole was a star of Forever Strong, but E2 shows he voiced Harvey Birdman, not a character in Forever Strong. Therefore, the claim is refuted. Status: Refuted.

### v2

@@ 3.1 Task Formulation
[…]
@@ 3.3.2 Reasoning Chain Generation and Selection
[…]
@@
[…]
@@ A.4 STRIVE-Generated Reasoning Chains
Claim: The star of Forever Strong played a character that is based on the Hanna-Barbera show “Birdman and the Galaxy Trio”. Label: Supported Model Output: C1: The star of Forever Strong played a character that is based on the Hanna-Barbara show “Birdman and the Galaxy Trio”. Entity Resolution: Star of Forever Strong -> Gary Cole (from E1) Resolution Verification: E1 confirms Gary Cole was a star of Forever Strong. Entity Resolution: Character played by Gary Cole -> ? (from E2)
Resolution Verification: E2 confirms Gary Cole voiced Harvey Birdman, but it doesn’t mention him playing a character in Forever Strong.
Verification: E1 shows Gary Cole was a star of Forever Strong, but E2 shows he voiced Harvey Birdman, not a character in Forever Strong. Therefore, the claim is refuted. Status: Refuted.